The volume of a room is 1,258 and three fifths cubic meters. The height of the room is 11 and three fifths meters. What is the area of the floor?
Found 3 solutions by Cromlix, Alan3354, checkley79:
Answer by Cromlix(4381)

The volume of a room is given by
length x breadth x height
For this room this equals 1,258 and 3/5 m^3
The area of the floor is given by
length x breadth.
Divide the volume of the room by the
height and you are left with the area of the
floor.
1,258.6 metres^3/ 11.6 metres = 108 1/2 metres^2

Answer by Alan3354(69443)

The volume of a room is 1,258 and three fifths cubic meters. The height of the room is 11 and three fifths meters. What is the area of the floor?
-----------------
Area = 1258.6/11.6
= 108.5 sq meters

Answer by checkley79(3341)

(1,258 3/5)/11 3/5=
[(1,258*5+3)/5]/[(11*5+3)/5]
[(6,290+3)/5]/[(55+3)]/5
[6,293/5]/[58/5] invert & multiply.
[6,293/5]